OpenStudy (help_people):

Two functions are shown in the table below. Function 1 2 3 4 5 6 f(x) = −x2 + 4x + 12 g(x) = x + 2 Complete the table on your own paper, then select the value that is a solution to f(x) = g(x). x = 2 x = 3 x = 5 x = 6

OpenStudy (help_people):

@triciaal @Elsa123

OpenStudy (help_people):

Hi :) no offense but are you almost done replying ?

OpenStudy (lynfran):

ok u need to input the given values into the function eq. f(x)=-x^2+4x+12 so f(1)=-(1)^2+4(1)+12 ... g(x)=x+2 then g(1)=(1)+2 ...

OpenStudy (help_people):

f(2)=-2^2+4(2)+12 g(x)=2+2

OpenStudy (help_people):

f(3)=-2^3+4(3)+12 g(x)=3+2

OpenStudy (help_people):

f(4)=-2^4+4(4)+12 g(x)=4+2

OpenStudy (lynfran):

u need to solve these...

OpenStudy (help_people):

could you at least help me :(?

OpenStudy (lynfran):

ok f(1)=-(1)^2+4(1)+12 =-1+4+12 =15 g(1)=(1)+2 =3

OpenStudy (help_people):

f(2)=-4+8+12 f(2)=4+12 f(2)=12 g(2)=4

OpenStudy (lynfran):

f(2)=-(2)^2+4(2)+12 =-(4)+8+12 =16 g(2)= (2)+2 =4

OpenStudy (lynfran):

f(3)=-(3)^2+4(3)+12 =-9+12+12 =15 g(3)=3+2 =5

OpenStudy (help_people):

i did f(2) wrong so are we done knee what o i ahebv to do next?

OpenStudy (lynfran):

f(4)=-(4)^2+4(4)+12 =-16+16+12 =12 g(4)= 4+2 =6 f(5)= -(5)^2+4(5)+12 = -25+20+12 =7 g(5)= 5+2 =7 f(6)= -(6)^2+4(6)+12 =-36+24+12 =0 g(6)=6+2 =8

OpenStudy (lynfran):

f(x)=g(x) -x^2+4x+12=x+2 -x^2+4x-x+12-2=0 -x^2+3x+10=0 -(x+2)(x-5)=0 x+2=0 x=-2 and x-5=0 x=5 now which option is it?

OpenStudy (lynfran):

did u realize that f(5) and g(5) are equal?

OpenStudy (help_people):

yes so x=5 is the option?

OpenStudy (lynfran):

correct

OpenStudy (help_people):

ty! could you help with one more??

OpenStudy (help_people):

actually a couple tbh???